# Randomly sampling bipartite networks with fixed degree sequences

### v3

@@ Abstract
Statistical analysis of bipartite networks frequently requires randomly sampling from the set of all bipartite networks with the same degree sequence as an observed network. Trade algorithms offer an efficient way to generate samples of bipartite networks by incrementally ‘trading’ the positions of some of their edges. However, it is difficult to know how many such trades are required to ensure that the sample is random. I propose a stopping rule that focuses on the distance between sampled networks and the observed network, and stops performing trades when this distribution stabilizes. Analyses demonstrate that, for over 300 different degree sequences, using this stopping rule ensures a random sample with a high probability, and that it is practical for use in empirical applications.
@@
journal: Submitted to arXiv
@@ 1 Introduction
In this paper, I propose a stopping rule for these algorithms that focuses on the distribution of sampled networks’ distances from the observed network. Using this rule, trades are performed until this distribution has stabilized according to a Kolmogorov–Smirnov (KS) test of distribution equality [Kolmogorov, 1933, Smirnov, 1948]. Examining this stopping rule’s performance in over 300 degree sequences, I demonstrate that it yields a random sample over 90% of the time.
@@ 2.1 Sampling bipartite networks using trade randomization
For any pair of degree sequences, there exists a (possibly empty) set $\mathcal{B}$ of all bipartite networks that have the given degree sequence. For example, given a top node degree sequence {1,1,2} and a bottom node degree sequence {1,1,2}, $\mathcal{B}$ contains the five networks shown in Figure 1A, and therefore the size or cardinality of $\mathcal{B}$ is five (i.e., $|\mathcal{B}|=5$). The goal of randomly sampling bipartite networks with fixed degree sequences is to sample networks from $\mathcal{B}$ such that each member of $\mathcal{B}$ has an equal probability of being sampled. The degree sequences {1,1,2} and {1,1,2} define a set that is small enough that it is possible to enumerate all its members, making it easy to directly sample from this list. However, because longer degree sequences define sets that are too large to enumerate, and indeed whose sizes are generally unknown [Barvinok, 2010], random sampling in such cases requires a different approach.
Many different approaches for sampling have been proposed [Verhelst, 2008, Miller & Harrison, 2013, Admiraal & Handcock, 2008, Bezáková et al., 2007, Chen et al., 2006, Fout et al., 2020, Wang, 2020, Strona et al., 2014, Godard & Neal, 2022], however they vary in the evidence of their randomness, their computational complexity, and their practicality in empirical contexts. The most widely-used sampling approach involves starting with a network (typically, the observed network being analyzed) that has the given degree sequences, then repeatedly rewiring its edges in a way that preserves the degree sequence, to obtain a new network that can be regarded as randomly sampled from $\mathcal{B}$. Among the most efficient algorithms proven to perform such rewiring randomly are the ‘trade’ algorithms [Strona et al., 2014, Godard & Neal, 2022, Carstens, 2015], so-called because they trade edges in the same way that children might trade baseball cards.
@@
Figure 2 illustrates the composition of a sample of 1000 networks drawn using the fastball trade algorithm, starting from Network B. Performing only one trade is insufficient to obtain a random sample of $\mathcal{B}$ because, as Figure 2A shows, it yields a sample in which Network B is over-represented and Network C is under-represented. The statistically significant $\chi^{2}$ statistic confirms that there is significant variation in the relative frequency of each network in the sample. This variation remains even after two (Figure 2B) or five (Figure 2C) trades have been performed. However, after ten trades have been performed (Figure 2D), each network occurs roughly the same number of times. The uniformity of the sample is confirmed by a non-significant $\chi^{2}$ statistic, which indicates that the sample is now a random sample from $\mathcal{B}$.
Carstens [2015] proved that trade algorithms will always yield a random sample from $\mathcal{B}$, provided a sufficient number of trades are performed. However, although trade algorithms are known to yield random samples, the number of trades required to ensure random sampling (i.e., the mixing time) is unknown and “intractable to compute” (p. 5). Therefore, in the next section, I propose a stopping rule to decide when a sufficient number of trades have been performed.
@@ 2.2 A trade randomization stopping rule
As Figure 2 illustrates, when $|\mathcal{B}|$ is small, a $\chi^{2}$ test can be used to determine when enough trades have been performed to obtain a random sample from $\mathcal{B}$. However, typically $|\mathcal{B}|$ is too large to enumerate all its members, and the $\chi^{2}$ test cannot be used to verify a sample’s randomness or to know when to stop performing trades. Therefore, a different method is required for deciding when a sufficient number of trades have been performed to obtain a random sample.
One alternative focuses not on the frequency of each network in the sample, but on the distance between each sampled network and the starting network. The distance $d$ between two bipartite networks is defined as the fraction of dyads that are different, or when the networks are represented as binary incidence matrices, the fraction of cells that are different [Strona et al., 2014]. For example, figure 1 illustrates that the distance between the starting network (Network B) and the ending network (Network A) is $d_{A,B}=0.444$.
Given a sample of bipartite networks obtained by randomizing a starting network using a trade algorithm, $D_{t}$ is the list of distances $d$ between the starting network and each network in the sample after performing $t$ trades. Figure 3 illustrates the distribution of $D$ in each sample whose composition is shown in figure 2. For example, figure 3A shows that, after one trade, the sample contains many networks at a distance of 0 from the starting network (these are Network B), some networks at a distance of 0.444 (these are Networks A, D, and E), and no networks at a distance of 0.666 (these would have been Network C). Note that this pattern mirrors the frequency of networks in the sample shown in figure 2A: many instances of Network B, some instances of Networks A, D, and E, and no instances of Network E.
Figure 3: Distribution of distance between a starting matrix and networks in a sample ($D$) drawn using a trade algorithm, following the given number of trades, with a Kolmogorov–Smirnov (KS) test of equality with an earlier distribution.
@@
Typically $\mathcal{D}$ is unknown, so it is not possible to directly test whether the distribution of $D$ matches the distribution of $\mathcal{D}$. However, because $D$ converges on $\mathcal{D}$, it is only necessary to test when the distribution of $D$ has stabilized. A Kolmogorov–Smirnov (KS) test of the equality of distributions [Kolmogorov, 1933, Smirnov, 1948] provides such a test when it is used to compare the distributions of $D_{t}$ and $D_{t-n}$. For example, the KS test reported in figure 3B ($KS=0.175,p<0.001$) shows that the distribution of $D$ after two trades is statistically significantly different from the distribution of $D$ after one trade; the distribution has not stabilized. In contrast, the KS test reported in figure 3D ($KS=0.023,p=0.954$) shows that the distribution of $D$ after ten trades is not statistically significantly different from the distribution of $D$ after five trade; after ten trades the distribution has stabilized. The stability of the distribution of $D_{1}0$ suggests that the sample is random after 10 trades, which was confirmed directly using a $\chi^{2}$ test in figure 2D.
@@ 3.1 Validation
The small example above illustrates one set of degree sequences (i.e., {1,1,2}, {1,1,2}) where, when the distribution of $D_{t}$ became not statistically significantly different from the distribution of $D_{t-n}$ (see figure 3D), the sample of networks was a random sample from $\mathcal{B}$ (see figure 2D). I follow a similar approach to test the hypothesis that the proposed stopping rule yields random samples, evaluating how often it yields a random sample when used to sample bipartite networks with other degree sequences.
For each degree sequence, I apply the stopping rule to draw a sample of 1000 bipartite networks, then evaluate whether the resulting sample is a random sample from $\mathcal{B}$. I repeat this 1000 times, recording the percentage of samples that are random. Applying the stopping rule requires computing $D$ and performing a KS test after every $n$ trades. In these analyses, to avoid needing to perform these computations too often, $n$ is twice the number of top nodes. The statistical significance of the KS test is evaluated using the conventional $p<0.05$ threshold.
To determine whether a given sample is random sample from $\mathcal{B}$, I require that it exhibit two properties: coverage and uniformity. First, the sample must include at least one instance of every member of, and therefore completely cover, $\mathcal{B}$. Second, the relative frequency of each network in the sample must be equal, and therefore be uniformly distributed, as indicated by a non-significant $chi^{2}$ test.
Using this method for determining whether a sample is a random sample from $\mathbf{B}$ requires that $|\mathcal{B}|$ is small enough for its members to be enumerated, and for a random sample of 1000 networks to plausibly contain at least one instance of each member. Therefore, I restrict the scope of this analysis to 335 different degree sequences where $|\mathcal{B}|<100$.
@@ 3.2 Application
The validation test evaluates whether the stopping rule consistently ensures a random sample, but it is only possible for shorter degree sequences that define smaller bipartite networks. To examine the stopping rule’s practicality in larger empirical bipartite networks, I illustrate its use to draw samples of bipartite networks in three different contexts: ecology, sociology, and political science.
@@
Each of these empirical examples involve a $\mathcal{B}$ that is too large to use a $chi^{2}$ test to determine whether a sample obtained using the stopping rule is random. Instead, in these illustrative applications, I focus on three outcomes that are observable. First, does the distribution of $D$ converge to a stable distribution that is expected to match the distribution of $\mathcal{D}$? Second, is each member of the resulting sample unique, suggesting that the sample provides broad coverage of $\mathcal{B}$ and does not over-represent certain members of $\mathcal{B}$. Third, can a likely-to-be-random sample be drawn in a practical amount of time using the stopping rule?
@@ 4.1 Validation
Table 1 reports, for 10 selected degree sequences and for the average over all 335 degree sequences, the percent of samples that were random when the stopping rule is used to stop performing trade randomizations. Each row represents a pair of bipartite network degree sequences, which define a $\mathcal{B}$ containing a specific number of networks ($|\mathcal{B}$). For example, the first row corresponds to the small example used throughout the background section. In this small example, using the proposed stopping rule to decide when to stop performing trade randomizations yielded a random sample of bipartite matrices from $\mathcal{B}$ 94.7% of the time. Using the stopping rule for other degree sequences yielded a random sample with similar frequency, and on average over 335 different degree sequences it yielded a random sample 93.2% of the time. This finding provides supports the hypothesis: the proposed stopping rule ensures with high probability that the resulting sample is a random sample from $\mathcal{B}$.
@@
{1,1,2} {1,1,2} 5 12 94.70%
{1,2,3} {1,1,2,2} 8 14 94.20%
{2,2,3} {1,1,1,2,2} 31 12 95.60%
{3,3,3} {1,1,1,2,2,2} 93 12 95.10%
{1,4,6} {1,1,1,2,2,2,2} 19 17 94.30%
{1,5,6} {1,1,1,1,2,2,2,2} 52 16 94.80%
{1,2,2,2} {1,1,2,3} 27 22 94.80%
{1,1,2,3} {1,1,1,2,2} 68 23 94.90%
{1,1,3,5} {1,1,1,2,2,3} 26 25 93.30%
{2,2,4,4,4} {2,2,4,4,4} 72 31 93.50%
Average for 335 sequences 43 23 93.20%
Results for all 335 degree sequences are provided in Supplementary Information.
@@ 4.2 Application
Figure 4 shows the distribution of $D$, after a specified number of trade randomizations, in a sample of bipartite networks with the same degree sequences as the (A) Darwin’s Finches, (B) Southern Women, and (C) US Senate networks. In each case, the stopping rule stopped performing trade randomizations when this distribution stabilized, which required a different number of trades depending on the network (e.g., 156 trades in the finches network, 1400 trades in the Senate network). It is impossible to verify that the resulting samples are random samples from the corresponding $\mathcal{B}$. However, it is possible to examine whether the stopping rule performs as expected and in practical time.
@@
Second, as required of a random sample, all 10,000 members of these samples are unique. This suggests that applying the stopping rule yields a sample with broad coverage of $\mathcal{B}$, and does not yield a sample in which any networks from $\mathcal{B}$ are over-represented.
Finally, using the stopping rule yielded a random sample of 10,000 bipartite networks in a practical amount of time: less then 4 seconds for the Finches and Women, and about 600 seconds (10 minutes) for the Senate. This is roughly 3.5 times longer than performing the same number of trades without periodically computing $D$ or a KS test (i.e., without applying the stopping rule). Although the additional computations required to use the stopping rule require more time, they offer confidence in the resulting sample’s randomness. For example, although 1400 trades can be performed on 10,000 copies of the Senate network in just 175 seconds, the resulting sample would come with no certainty about its randomness because more (or fewer) trades may have been required.

### v4

@@ Abstract
Statistical analysis of bipartite networks frequently requires randomly sampling from the set of all bipartite networks with the same degree sequence as an observed network. Trade algorithms offer an efficient way to generate samples of bipartite networks by incrementally ‘trading’ the positions of some of their edges. However, it is difficult to know how many such trades are required to ensure that the sample is random. I propose a stopping rule that focuses on the distance between sampled networks and the observed network, and stops performing trades when this distribution stabilizes. Analyses demonstrate that, for over 650 different degree sequences, using this stopping rule ensures a random sample with a high probability, and that it is practical for use in empirical applications.
@@
journal: arXiv pre-print
@@ 1 Introduction
In this paper, I propose a stopping rule for these algorithms that focuses on the distribution of sampled networks’ distances from the observed network. Using this rule, trades are performed until this distribution has stabilized. Examining this stopping rule’s performance in over 650 degree sequences, I demonstrate that on average it yields a random sample over 85% of the time.
@@ 2.1 Sampling bipartite networks using trade randomization
For any pair of degree sequences, there exists a (possibly empty) set $\mathcal{B}$ of all bipartite networks that have the given degree sequence. For example, given a top node degree sequence {1,2,1} and a bottom node degree sequence {1,2,1}, $\mathcal{B}$ contains the five networks shown in Figure 1A, and therefore the size or cardinality of $\mathcal{B}$ is five (i.e., $|\mathcal{B}|=5$). The goal of randomly sampling bipartite networks with fixed degree sequences is to sample networks from $\mathcal{B}$ such that each member of $\mathcal{B}$ has an equal probability of being sampled. The degree sequences {1,2,1} and {1,2,1} define a set that is small enough that it is possible to enumerate all its members, making it easy to directly sample from this list. However, because longer degree sequences define sets that are too large to enumerate, and indeed whose sizes are generally unknown [Barvinok, 2010], random sampling in such cases requires a different approach.
Many different approaches for sampling have been proposed [Verhelst, 2008, Miller & Harrison, 2013, Admiraal & Handcock, 2008, Bezáková et al., 2007, Chen et al., 2006, Fout et al., 2020, Wang, 2020, Strona et al., 2014, Godard & Neal, 2022], however they vary in the evidence of their randomness, their computational complexity, and their practicality in empirical contexts. The most widely used sampling approach involves starting with a network (typically, the observed network being analyzed) that has the given degree sequences, then repeatedly rewiring its edges in a way that preserves the degree sequence, to obtain a new network that can be regarded as randomly sampled from $\mathcal{B}$. Among the most efficient algorithms proven to perform such rewiring randomly are the ‘trade’ algorithms [Strona et al., 2014, Godard & Neal, 2022, Carstens, 2015], so-called because they trade edges in the same way that children might trade baseball cards.
@@
Figure 2 illustrates the composition of a sample of 1000 networks drawn using the fastball trade algorithm, starting from Network B. Performing only one trade is insufficient to obtain a random sample from $\mathcal{B}$ because, as Figure 2A shows, it yields a sample in which Network B is over-represented and Network C is under-represented. The statistically significant $\chi^{2}$ statistic confirms that there is significant variation in the relative frequency of each network in the sample. This variation remains even after two (Figure 2B) or five (Figure 2C) trades have been performed. However, after ten trades have been performed (Figure 2D), each network occurs roughly the same number of times. The uniformity of the sample is confirmed by a non-significant one-sample $\chi^{2}$ statistic, which indicates that the sample is now a random sample from $\mathcal{B}$.
Prior work on trade algorithms has focused on proving that they yield random samples from $\mathcal{B}$ if a sufficient number of trades are performed [Carstens, 2015], or on developing ways to perform those trades more quickly [e.g., Godard & Neal, 2022, Carstens & Kleer, 2018]. However, no prior work has directly investigated how many such trades are required to ensure random sampling (i.e., the mixing time), which is unknown and “intractable to compute” [Carstens, 2015, p. 5]. When Strona et al. [2014] initially developed the curveball trade algorithm, they provided implementations of their algorithm in R and Python, in which they fixed the number of trades performed at $5T$, where $T$ is the number of top nodes. However, they did not provide a rationale for choosing this value, or evidence supporting its adequacy for yielding random samples. Therefore, in the next section, I propose a stopping rule to decide when a sufficient number of trades have been performed.
@@ 2.2 A trade randomization stopping rule
As Figure 2 illustrates, when $|\mathcal{B}|$ is small, a one-sample $\chi^{2}$ test can be used to determine when enough trades have been performed to obtain a random sample from $\mathcal{B}$. However, typically $|\mathcal{B}|$ is too large to enumerate all its members, and the $\chi^{2}$ test cannot be used to verify a sample’s randomness or to know when to stop performing trades. Therefore, a different method is required for deciding when a sufficient number of trades have been performed to obtain a random sample.
One alternative focuses not on the frequency of each network in the sample, but on the distance between each sampled network and the starting network. The distance $d$ between two bipartite networks is defined as the fraction of dyads that are different, or when the networks are represented as binary incidence matrices, the fraction of cells that are different [Strona et al., 2014]. Formally
$d_{A,B}=\frac{\sum{|A-B|}}{r\times c},$
where $A$ and $B$ are binary $r\times c$ incidence matrices. For example, Figure 1 illustrates that the distance between the starting network (Network B) and the ending network (Network A) is $d_{A,B}=0.444$.
Given a sample of bipartite networks obtained by randomizing a starting network using a trade algorithm, $D_{t}$ is the list of distances $d$ between the starting network and each network in the sample after performing $t$ trades. Figure 3 illustrates the distribution of $D$ in each sample whose composition is shown in Figure 2. For example, Figure 3A shows that, after one trade, the sample contains many networks at a distance of 0 from the starting network (these are Network B), some networks at a distance of 0.444 (these are Networks A, D, and E), and no networks at a distance of 0.666 (these would have been Network C). Note that this pattern mirrors the frequency of networks in the sample shown in Figure 2A: many instances of Network B, some instances of Networks A, D, and E, and no instances of Network C.
Figure 3: Distribution of distance between a starting matrix and networks in a sample ($D$) drawn using a trade algorithm, following the given number of trades, with an Anderson-Darling (AD) test of equality with an earlier distribution.
@@
Typically $\mathcal{D}$ is unknown, so it is not possible to directly test whether the distribution of $D$ matches the distribution of $\mathcal{D}$. However, because $D$ converges on $\mathcal{D}$, it is only necessary to test when the distribution of $D$ has stabilized. There are many ways to test whether $D_{t}$ and $D_{t-1}$ were drawn from the same distribution. The Kolmogorov–Smirnov test [Kolmogorov, 1933, Smirnov, 1948] is the most widely known such test, however because lacks sensitivity in a range of cases, two other tests are more appropriate in this context. When the values in $D$ are unequally spaced (i.e., ordinal; as they are in the example illustrated in Figure 3), the Anderson-Darling (AD) test [Anderson & Darling, 1952] is optimal [Dowd, 2020]. In contrast, when the values in $D$ are equally spaced (i.e., interval), the DTS test is optimal [Dowd, 2020].
The AD test reported in Figure 3B shows that the distribution of $D$ after two trades is statistically significantly different from the distribution of $D$ after one trade; the distribution has not stabilized. In contrast, the AD test reported in Figure 3D shows that the distribution of $D$ after ten trades is not statistically significantly different from the distribution of $D$ after five trades; after ten trades the distribution has stabilized. The stability of the distribution of $D_{10}$ suggests that the sample is random after 10 trades, which was confirmed directly by the $\chi^{2}$ test reported in Figure 2D.
@@ 3.1 Validation
The small example above illustrates one set of degree sequences (i.e., {1,2,1}, {1,2,1}) where, when the distribution of $D_{t}$ became not statistically significantly different from the distribution of $D_{t-n}$ (see Figure 3D), the sample of networks was a random sample from $\mathcal{B}$ (see Figure 2D). I follow a similar approach to test the hypothesis that the proposed stopping rule yields random samples, evaluating how often it yields a random sample when used to sample bipartite networks with other degree sequences. Specifically, I consider 665 different degree sequences that define $\mathcal{B}$ with $|\mathcal{B}|\leq 1000$.
For each degree sequence, I apply the stopping rule to draw a sample of 1000 bipartite networks, then evaluate whether the resulting sample is a random sample from $\mathcal{B}$. I repeat this 100 times, recording the percentage of samples that are random. Applying the stopping rule requires computing $D$ and performing a AD or DTS test after every $n$ trades. To avoid needing to perform these computations too often, $n$ is equal to the number of top nodes. The statistical significance of the test is evaluated using the conventional $p<0.05$ threshold.
To determine whether a given sample is random sample from $\mathcal{B}$, following the approach illustrated in Figure 2, I use a one-sample $\chi^{2}$ test. A sample is uniformly random when this test is not statistically significant, indicating that the distribution of matrix frequencies does not differ from a uniform distribution.
@@ 3.2 Application
The validation test evaluates whether the stopping rule consistently ensures a random sample, but it is only possible for shorter degree sequences that define $\mathcal{B}$ with smaller $|\mathcal{B}|$. To examine the stopping rule’s practicality for randomly sampling from larger spaces, I illustrate its use to draw samples of bipartite networks in three different contexts: ecology, sociology, and political science. In each case, I draw a sample of 1000 random bipartite networks because this is the typical size of a sample required to evaluate the significance of patterns in the network of interest [Strona et al., 2014].
@@
Each of these empirical examples involve a $\mathcal{B}$ that is too large to use a $\chi^{2}$ test to determine whether a sample obtained using the stopping rule is random. Instead, in these illustrative applications, I focus on three outcomes that are observable. First, does the distribution of $D$ converge to a stable distribution that is expected to match the distribution of $\mathcal{D}$? Second, is each member of the resulting sample unique, suggesting that the sample provides broad coverage of $\mathcal{B}$ and does not over-represent certain members of $\mathcal{B}$. Third, can a likely-to-be-random sample be drawn in a practical amount of time using the stopping rule?
@@ 4.1 Validation
Table 1 reports, for 10 selected degree sequences and for the average over all 665 degree sequences, the percent of samples that were random when the stopping rule is used to decide when to stop performing trade randomizations. Each row represents a pair of bipartite network degree sequences, which define a $\mathcal{B}$ containing a specific number of networks ($|\mathcal{B}|$). For example, the first row corresponds to the small example used throughout the background section. In this small example, using the proposed stopping rule to decide when to stop performing trade randomizations yielded a random sample of bipartite matrices from $\mathcal{B}$ 92% of the time. Using the stopping rule for other degree sequences yielded a random sample with similar frequency, and on average over 365 different degree sequences it yielded a random sample 87.5% of the time.
@@
{1,2,2} {1,2,2} 5 10 92%
{2,2,3} {1,2,2,2} 12 10 95%
{1,3,3} {1,1,1,2,2} 18 10 88%
{1,4,4} {1,1,1,2,2,2} 24 11 91%
{1,2,5} {1,1,1,1,1,1,2} 51 11 90%
{1,3,6} {1,1,1,1,1,1,2,2} 72 11 93%
{2,2,2,3} {1,2,3,3} 27 15 92%
{2,2,2,2} {1,1,1,2,3} 108 14 96%
{1,1,3,4} {1,1,1,1,2,3} 54 18 89%
{1,2,3,4,4} {1,3,3,3,4} 76 24 84%
Average for 665 sequences 172 17 87.50%
Results for all 665 degree sequences are provided in Supplementary Information.
These estimated rates of achieving random samples must be interpreted with caution because they are likely underestimated. The one-sample $\chi^{2}$ test used to determine whether a sample is uniformly random is performed with $\alpha=0.05$, which gives the test a type-I error rate of 5%. This means that the test will incorrectly reject the null hypothesis (i.e., find that the sample is not random) despite the null hypothesis being correct (i.e., the sample actually is random) in 5% of cases. Therefore, even if the stopping rule yielded a random sample 100% of the time, the $\chi^{2}$ test would find it did so only 95% of the time. Adjusting these rates to account for this type-I error suggests that the stopping rule’s average success rate is 92% (i.e., $\frac{0.875}{0.95}$).
Whether the raw or adjusted success rate is used, these findings provide supports the hypothesis: the proposed stopping rule ensures with high probability that the resulting sample is a random sample from $\mathcal{B}$.
@@ 4.2 Application
Figure 4 shows the distribution of $D$, after a specified number of trade randomizations, in a sample of bipartite networks with the same degree sequences as the (A) Darwin’s Finches, (B) Southern Women, and (C) US Senate networks. In each case, the stopping rule stopped performing trade randomizations when this distribution stabilized, which required a different number of trades depending on the network (e.g., 117 trades in the finches network, 1100 trades in the Senate network). It is impossible to verify that the resulting samples are random samples from the corresponding $\mathcal{B}$. However, it is possible to examine whether the stopping rule performs as expected and in practical time.
@@
Second, as required of a random sample, all 1,000 members of these samples are unique. This suggests that applying the stopping rule yields a sample with broad coverage of $\mathcal{B}$, and does not yield a sample in which any networks from $\mathcal{B}$ are over-represented.
Finally, using the stopping rule yielded a random sample of 1,000 bipartite networks in a practical amount of time: about 1 second for the Finches and Women, and about 77 seconds for the Senate, using an Apple M1 Max processor with 64GB of RAM.